The on-line phone book Case for its use as a University resource

Douglas Rogers - Department of Computer Science

November 24, 1994

The on-line phone book is a personal information database accessible through the Internet. It uses the programs written by the Computer and Communications Services Office at the University of Illinois at Urbana. It is now used at a number of American universities primarily for the normally available information on students. It is based on a flat database representation and a simple querying mechanism.

It provides mechanisms to define mail aliases and can be interrogated by any device with Ethernet access. There are a number of user interface tools - of which there are good ones for both PC’s and MAC’s, though for Unix systems the command line interface is good, but the X viewer appears broken. There is also a viewer for the web.

Being query-able across the Internet it is straightforward to write perl scripts to extract the required information for other purposes. There is also comment in the FAQ for the ph-lists newsgroup, about mail re-direction tools.

Why use PH in Edinburgh?

• Because it is already written

• Because it is accessible across all architectures

• Because the code will be maintained in future

• Because it is easy to put up maintain and support

• Because it doesn’t provide just aliases, but acts as a phone book

• Because it will also allow us to set up user accounts with no paperwork and no re-entering of data.

So what are the problems?

• Politics - without an agreed set of fields and how they will be accessed and used then there is no point in starting.

• A good x windows interface needs finding or writing.

• Ensuring that the new data on students is fed in as early as possible

• Getting all the data on all the staff together

• Persuading people that it is a good idea

• Issuing passwords to everyone

• Providing a support mechanism